# Continuity of asymptotic entropy on wreath products

### v1

@@ 1.5. Acknowledgements
I would like to thank Anna Erschler for suggesting this project, for indicating the reference [Cou96] to me, and for many helpful discussions as well as comments on the first draft of this paper. I am grateful to Inhyeok Choi for indicating the reference [AAV13] for the proof of Lemma 3.5, as well as for helpful comments and references for the applications of Theorem 1.5 to acylindrically hyperbolic groups. I also thank Petr Kosenko for many thought-provoking discussions around the continuity of harmonic measures on the Poisson boundary. I am very grateful to Anna Cascioli for her careful reading of the first version of this manuscript. The author is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2044 –390685587, Mathematics Münster: Dynamics–Geometry–Structure.
@@ 2.1. Random walks on groups, entropy and the Poisson boundary
The space $(G^{\mathbb{Z}_{+}},\mathbb{P}_{\mu})$ is called the space of sample paths or the space of trajectories of the $\mu$-random walk. When there is no risk of confusion, we will denote $\mathbb{P}=\mathbb{P}_{\mu}$. However, in most of the paper we will be working at the same time with a probability measure $\mu$ as well as with a sequence of probability measures $\{\mu_{k}\}_{k\geq 1}$. In this context we will need to distinguish between $\mathbb{P}_{\mu}$ and $\mathbb{P}_{\mu_{k}}$, $k\geq 1$.
@@
For every $n\geq 1$, define the partition $\alpha_{n}$ of the space of sample paths $G^{\mathbb{N}}$, where two trajectories $\mathbf{w},\mathbf{w^{\prime}}\in G^{\mathbb{Z}_{+}}$ belong to the same element of $\alpha_{n}$ if and only if $w_{n}=w^{\prime}_{n}$.
@@
Recall that a countable group $G$ is called ICC if every non-trivial element in the associated quotient group has an infinite conjugacy class. The hyper-FC-center of a group $G$ is the minimal normal subgroup of $G$ with the property that the associated quotient group is ICC. A group $G$ is said to be hyper-FC-central if it coincides with its hyper-FC-center. A finitely generated group is hyper-FC central if and only if it contains a nilpotent subgroup of finite index [McL56, Theorem 2]. It is known that a countable group $G$ is hyper-FC-central if and only if every probability measure on $G$ has the Liouville property. The fact that hyper-FC-central groups have this property goes back to [Bla55, CD60, DSW60] for abelian groups, to [DM61, Mar66] for nilpotent groups, and to [LZ98, Jaw04] for the general case. Reciprocally, it is proved in [FHTVF19, Theorem 1] that any countable group that is not hyper-FC-central admits a non-degenerate probability measure with finite entropy and without the Liouville property.
@@
Consider countable partitions $\rho,\gamma$ and $\delta$ of $G^{\mathbb{Z}_{+}}$. The following properties hold:
@@ 2.2. Escape probability and the asymptotic range
$\lim_{n\to\infty}\frac{R_{n}}{n}=p_{\mathrm{esc}}(\mu)=\mathbb{P}(w_{n}\neq e_{G}\text{ for all }n\geq 1).$ (2)
@@ 2.4. Growth of groups
We finish this section by recalling the basic notions related to the growth of finitely generated groups. We refer to [dlH00, Chapter VI] for a more detailed exposition.
@@ 3. Convergence of probability measures
The right hand side converges to 00 as $k\to\infty$ since we have convergence in total variation, and hence we conclude the proof.
@@
$\mathbb{P}_{\mu_{k}}\left(\exists\ 1\leq\ell\leq n\text{ s.t. }w_{\ell}=g\right)\xrightarrow[k\to\infty]{}\mathbb{P}_{\mu}\left(\exists\ 1\leq\ell\leq n\text{ s.t. }w_{\ell}=g\right).$
@@
The following proposition will be used several times in Section 5. Recall from Definition 1.3 that $p_{\mathrm{esc}}(\mu)\coloneqq\mathbb{P}(w_{n}\neq e_{G}\text{ for every }n\geq 1)$.
@@
$\mathbb{P}_{\mu_{k}}\left(\exists\ 1\leq\ell\leq n_{0}\text{ such that }w_{\ell}=e_{G}\right)\xrightarrow[k\to\infty]{}\mathbb{P}_{\mu}\left(\exists\ 1\leq\ell\leq n_{0}\text{ such that }w_{\ell}=e_{G}\right).$ (3)
@@
$\mathbb{P}_{\mu_{k}}\left(\exists\ell>n_{0}\text{ such that }w_{\ell}=e_{G}\right)\xrightarrow[k\to\infty]{}\mathbb{P}_{\mu}\left(\exists\ell>n_{0}\text{such that }w_{\ell}=e_{G}\right)$ (4)
@@ 4.1. The comparison lemma of Coulhon and Saloff-Coste
    Let $\check{\mu}$ be the reflected measure of $\mu$, i.e. $\check{\mu}(g)=\mu(g^{-1})$ for each $g\in G$. Consider $1\leq p\leq q\leq\infty$, and $p^{\prime},q^{\prime}$ such that $11/p+1/p^{\prime}=1/q+1/q^{\prime}=$ 1. Then the adjoint operator of $T_{\mu}:\ell^{p}(G)\to\ell^{q}(G)$ is $T_{\check{\mu}}:\ell^{q^{\prime}}(G)\to\ell^{p^{\prime}}(G)$. In particular, $\|T_{\mu}\|_{p\to q}=\|T_{\check{\mu}}\|_{q^{\prime}\to p^{\prime}}$.
The following proposition is proved in [Cou96, Proposition IV.4] (see also [CSC90, Section II]), without keeping track of what is the precise value of the constant $C$. The proof that we present here is a self-contained recollection from the results of [Cou96, Section IV], with minor modifications to control the value of $C$ that appears in the statement of the proposition.
@@
As it is shown in the proof of [Cou96, Proposition IV.2], from Equation (5) and the assumption that $\mu_{1}$ is symmetric it follows that for each $f\in\ell^{1}(G)\cap\ell^{2}(G)\backslash\{0\}$ with $\|f\|_{1}\leq 1$, we have
@@
Let $f\in\ell^{1}(G)\cap\ell^{2}(G)\backslash\{0\}$ with $\|f\|_{1}\leq 1$, and recall that Equation (6) is satisfied. Let us consider the following three cases, which together prove that Equation (12) below holds.
@@
Note that $\xi$ and the sequence $\{\tilde{m}_{n}\}_{n\geq 1}$ are defined only in terms of the constants $C_{1}$ and $C_{2}$. Recall that at the beginning of the proof we defined $m(x)\coloneqq C_{2}x^{-d/2}$. Next, by following the proof of [Cou96, Theorem IV.3], we see that there is $K\geq 1$ such that $\widetilde{m}_{Kn}\leq m(n)$, for all $n\geq 1$. Indeed, the sequence $\{\tilde{m}_{n}\}_{n\geq 1}$ is decreasing and converges to 00 for $n\to\infty$. Therefore, we can choose $K\geq 1$ sufficiently large such that
@@
Since the growth function of the group $H\coloneqq\langle\mathrm{supp}(\mu)\rangle$ satisfies $v_{H}(n)\succcurlyeq n^{d}$ and $\nu_{S}$ is symmetric and finitely supported, it is a classical result of Varopoulos [Var87] (see also [VSCC92, Theorems VI.3.3 & VI.5.1]) that there exists a constant $C_{3}>0$ such that
@@
[…]
@@
[…]
@@
[…]
@@ 5.1. The coarse trajectory
That is, $\mathcal{P}_{n}^{t_{0}}(G)$ consists of an ordered collection of the group element visited by the random walk, every $t_{0}$ steps, between time 00 and $n$.
@@
By repeating this argument $s-1$ times, we obtain the inequality
@@ 5.2. The partition of bad increments
We now partition the time interval between 00 and $n$ into subintervals of length $t_{0}$. Afterward, we define what it means for these intervals to be good or bad, depending on whether the increments of the random walk during the time instants belonging to the interval belong to a given fixed finite subset or not.
@@
The following is a version of the “obscuring lemma” which appears in [CFFT22, Lemma 2.4] and [FS23, Lemma 3.7], that is uniform along the sequence of probability measures $\mu_{k}$.
@@
For this proof, we introduce the notation $\kappa:[0,1]\to[0,1]$ for the function $\kappa(x)\coloneqq-x\log(x)$, where we use the standard convention $00\log(0)=$ 0. The function $\kappa$ is continuous, concave, strictly increasing in the interval $[0,1/e]$ and strictly decreasing in the interval $[1/e,1].$
@@
$H_{\mu_{k}}(\widetilde{W_{j}})$ $=\kappa(\mu^{t_{0}}_{k}(\Omega\backslash E))+\sum_{d\in Q}\kappa(\mu^{t_{0}}_{k}(\{d\}\cap E))+\sum_{d\in D\backslash Q}\kappa(\mu^{t_{0}}_{k}(\{d\}\cap E))$
@@
$\mu_{k}^{t_{0}}(I_{j}\text{ is an }(L,R)\text{-bad interval})\leq t_{0}\mu_{k}(\{g\in G\mid g\text{ is }(L,R)\text{-bad}\})<\delta\text{ for all }k\geq K_{3}.$
@@
[…]
@@ 6.1. Unstable elements and their visit times
[…]
@@
[…]
@@
[…]
@@
This is a consequence of the fact that if $m_{1}\geq m_{2}$, then the event where 00 is poorly $m_{1}$-stabilized is contained in the event where 00 is poorly $m_{2}$-stabilized. ∎
@@
[…]
@@
[…]
@@
Let $t_{0}\geq 1$ and consider any finite subset $F\subseteq B$ with $e_{B}\in F$ Then for every $\varepsilon>0$, there exists $K,n_{0}\geq 1$ such that for all $k\geq K$ and $n>n_{0}t_{0}$ we have $H_{\mu_{k}}(\mathcal{U}_{n}(n_{0},t_{0},F)\mid\mathcal{P}_{n}^{t_{0}})<\varepsilon n.$
@@
$>\displaystyle J_{m}\coloneqq\{j\in\mathbb{Z}_{\geq 0}\mid\text{ there exist }n_{1}>n_{2}>$$\ldots>n_{m}>j\text{ such that }$
@@
$\mathbb{P}_{\mu_{k}}(0\in J_{m})$ $\leq\mathbb{P}_{\mu_{k}}(FF^{-1}\text{ is visited at least }m\text{ times in the future})$
@@
$\mathbb{E}(|\mathcal{V}_{n}(n_{0},t_{0},F)\backslash J_{n_{0}}|)\leq n_{0}|F|\mathbb{E}_{k}(|\mathcal{U}_{n}(n_{0},t_{0},F)|).$
@@
$H_{\mu_{k}}\Big{(}\mathcal{V}_{n}(n_{0},t_{0},F))\leq\left(\frac{n}{t_{0}}-n_{0}+1\right)\log\left(2\right)\leq\left(\frac{n}{T}\right)\log\left(2\right)=\varepsilon,$

### v2

@@ 1.5. Acknowledgements
I would like to thank Anna Erschler for suggesting this project, for indicating the reference [COU96] to me, and for many helpful discussions as well as comments on the first draft of this paper. I am grateful to Inhyeok Choi for indicating the reference [AAV13] for the proof of Lemma 3.5, as well as for helpful comments and references for the applications of Theorem 1.5 to acylindrically hyperbolic groups. I also thank Petr Kosenko for many thought-provoking discussions around the continuity of harmonic measures on the Poisson boundary. I am very grateful to Anna Cascioli for her careful reading of the first version of this manuscript. I would like to thank the anonymous referee for their careful reading of this paper, as well as for suggesting the addition of Proposition 8.12 and Corollary 8.13. The author is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2044 –390685587, Mathematics Münster: Dynamics–Geometry–Structure.
@@ 2.1. Random walks on groups, entropy and the Poisson boundary
The space $(G^{\mathbb{Z}_{+}},\mathbb{P}_{\mu})$ is called the space of sample paths or the space of trajectories of the $\mu$-random walk. In most of the paper we will be working at the same time with a probability measure $\mu$ as well as with a sequence of probability measures $\{\mu_{k}\}_{k\geq 1}$. In this context we will need to distinguish between $\mathbb{P}_{\mu}$ and $\mathbb{P}_{\mu_{k}}$, $k\geq 1$.
@@
For every $n\geq 1$, define the partition $\alpha_{n}$ of the space of sample paths $G^{\mathbb{Z}}$, where two trajectories $\mathbf{w},\mathbf{w^{\prime}}\in G^{\mathbb{Z}_{+}}$ belong to the same element of $\alpha_{n}$ if and only if $w_{n}=w^{\prime}_{n}$.
@@
Recall that a countable group $G$ is called ICC if every non-trivial element in the associated quotient group has an infinite conjugacy class. The hyper-FC-center of a group $G$ is the minimal normal subgroup of $G$ with the property that the associated quotient group is ICC. A group $G$ is said to be hyper-FC-central if it coincides with its hyper-FC-center. A finitely generated group is hyper-FC central if and only if it contains a nilpotent subgroup of finite index [MCL56, Theorem 2]. It is known that a countable group $G$ is hyper-FC-central if and only if every probability measure on $G$ has the Liouville property. The fact that hyper-FC-central groups have this property goes back to [BLA55, CD60, DSW60] for abelian groups, to [DM61, MAR66] for nilpotent groups, and to [LZ98, JAW04] for the general case. Reciprocally, it is proved in [FHT+19, Theorem 1] that any countable group that is not hyper-FC-central admits a non-degenerate probability measure with finite entropy and without the Liouville property.
@@
Consider countable partitions $\rho,\gamma$ and $\delta$ of a Borel space. The following properties hold:
@@ 2.2. Escape probability and the asymptotic range
$\lim_{n\to\infty}\frac{R_{n}}{n}=p_{\mathrm{esc}}(\mu)=\mathbb{P}_{\mu}(w_{n}\neq e_{G}\text{ for all }n\geq 1).$ (2)
@@ 2.4. Growth of groups
We finish this section by recalling the basic notions related to the growth of finitely generated groups. We refer to [DE 00, Chapter VI] for a more detailed exposition.
@@ 3. Convergence of probability measures
The right hand side converges to 0 as $k\to\infty$ since we have convergence in total variation, and hence we conclude the proof.
@@
[…]
@@
The following proposition will be used several times in Section 5. Recall from Definition 1.3 that $p_{\mathrm{esc}}(\mu)\coloneqq\mathbb{P}_{\mu}(w_{n}\neq e_{G}\text{ for every }n\geq 1)$.
@@
[…]
@@
[…]
@@ 4.1. The comparison lemma of Coulhon and Saloff-Coste
    Let $\check{\mu}$ be the reflected measure of $\mu$, i.e. $\check{\mu}(g)=\mu(g^{-1})$ for each $g\in G$. Consider $1\leq p\leq q\leq\infty$, and $p^{\prime},q^{\prime}$ such that $1/p+1/p^{\prime}=1/q+1/q^{\prime}=1$. Then the adjoint operator of $T_{\mu}:\ell^{p}(G)\to\ell^{q}(G)$ is $T_{\check{\mu}}:\ell^{q^{\prime}}(G)\to\ell^{p^{\prime}}(G)$. In particular, $\|T_{\mu}\|_{p\to q}=\|T_{\check{\mu}}\|_{q^{\prime}\to p^{\prime}}$.
Proof.
Items (1), (2) and (3) can be checked by direct computation. For Item (4), we first note that with $p,p^{\prime},q,q^{\prime}$ as in the statement it holds that $\ell^{p^{\prime}}(G)$ is the dual of $\ell^{p}(G)$ and that $\ell^{q^{\prime}}(G)$ is the dual of $\ell^{q}(G)$. In addition, for any $f_{1}\in\ell^{p}(G)$ and $f_{2}\in\ell^{q}(G)$ we have
$\langle T_{\mu}f_{1},f_{2}\rangle$ $=\sum_{g\in G}\sum_{h\in G}f_{1}(gh)\mu(h)f_{2}(g)$
$=\sum_{h\in G}\sum_{g\in G}f_{1}(gh)\mu(h)f_{2}(g)$
$=\sum_{h\in G}\sum_{g\in G}f_{1}(g)\mu(h)f_{2}(gh^{-1})$
$=\sum_{g\in G}\sum_{h\in G}f_{1}(g)\check{\mu}(h)f_{2}(gh)$
$=\langle f_{1},T_{\check{\mu}}f_{2}\rangle.$
This shows that $T_{\check{\mu}}$ is the adjoint operator of $T_{\mu}$. The last equality in the statement of Item (4) follows from the general fact that the an operator between Banach spaces has the same norm as its adjoint. ∎
The following proposition is proved in [COU96, Proposition IV.4] (see also [CS90, Section II]), without keeping track of what is the precise value of the constant $C$. The proof that we present here is a self-contained recollection from the results of [COU96, Section IV], with minor modifications to control the value of $C$ that appears in the statement of the proposition.
@@
[…]
@@
[…]
@@
Note that $\xi$ and the sequence $\{\tilde{m}_{n}\}_{n\geq 1}$ are defined only in terms of the constants $C_{1}$ and $C_{2}$. Recall that at the beginning of the proof we defined $m(x)\coloneqq C_{2}x^{-d/2}$. Next, by following the proof of [COU96, Theorem IV.3], we see that there is $K\geq 1$ such that $\widetilde{m}_{Kn}\leq m(n)$, for all $n\geq 1$. Indeed, the sequence $\{\tilde{m}_{n}\}_{n\geq 1}$ is decreasing and converges to 0 for $n\to\infty$. Therefore, we can choose $K\geq 1$ sufficiently large such that
@@
Since the growth function of the group $H\coloneqq\langle\mathrm{supp}(\mu)\rangle$ satisfies $v_{H}(n)\succcurlyeq n^{d}$ and $\nu_{S}$ is symmetric and finitely supported, it is a classical result of Varopoulos [VAR87] (see also [VSC92, Theorems VI.3.3 & VI.5.1]) that there exists a constant $C_{3}>0$ such that
@@
$\frac{1}{p_{\mathrm{esc}}(\mu)}=\sum_{n=0}^{\infty}\mu^{*n}(e_{G})=\lim_{t\to 1^{-}}\frac{1}{(2\pi)^{d}}\int_{[-\pi,\pi]^{d}}\frac{1}{\mathrm{Re}\left(1-t\phi_{\mu}(\mathbf{k})\right)}\ d\mathbf{k};$
@@
$\frac{1}{p_{\mathrm{esc}}(\mu)}=\frac{1}{(2\pi)^{d}}\int_{[-\pi,\pi]^{d}}\frac{1}{\mathrm{Re}\left(1-\phi_{\mu}(\mathbf{k})\right)}\ d\mathbf{k}<\infty.$
@@
$\frac{(2\pi)^{d}}{p_{\mathrm{esc}}(\mu_{k})}=\int_{[-\pi,\pi]^{d}}\frac{1}{\mathrm{Re}\left(1-\phi_{\mu_{k}}(\mathbf{k})\right)}\ d\mathbf{k}\xrightarrow[k\to\infty]{}\int_{[-\pi,\pi]^{d}}\frac{1}{\mathrm{Re}\left(1-\phi_{\mu}(\mathbf{k})\right)}\ d\mathbf{k}=\frac{(2\pi)^{d}}{p_{\mathrm{esc}}(\mu)},$
@@ 5.1. The coarse trajectory
That is, $\mathcal{P}_{n}^{t_{0}}(G)$ consists of an ordered collection of the group element visited by the random walk, every $t_{0}$ steps, between time 0 and $n$.
@@
By repeating this argument s1s-1 times, we obtain the inequality
@@ 5.2. The partition of bad increments
We now partition the time interval between 0 and $n$ into subintervals of length $t_{0}$. Afterward, we define what it means for these intervals to be good or bad, depending on whether the increments of the random walk during the time instants belonging to the interval belong to a given fixed finite subset or not.
@@
The following is a version of the “obscuring lemma” which appears in [CFF+25, Lemma 2.4] and [FS23, Lemma 3.7], that is uniform along the sequence of probability measures $\mu_{k}$.
@@
For this proof, we introduce the notation $\kappa:[0,1]\to[0,1]$ for the function $\kappa(x)\coloneqq-x\log(x)$, where we use the standard convention $0\log(0)=0$. The function $\kappa$ is continuous, concave, strictly increasing in the interval $[0,1/e]$ and strictly decreasing in the interval $[1/e,1].$
@@
[…]
@@
[…]
@@
$E_{j}\coloneqq\left\{I_{j}\text{ is an }(L,R)\text{-bad interval}\right\}.$
@@ 6.1. Unstable elements and their visit times
$\mathbb{P}_{\mu_{k}}\left(j\text{ is poorly }m\text{-stabilized}\right)\leq\mathbb{P}_{\mu_{k}}\left(0\text{ is poorly }m\text{-stabilized}\right).$
@@
$\mathbb{P}_{\mu_{k}}(j\text{ is poorly }m\text{-stabilized})$ $=\mathbb{P}_{\mu_{k}}(\text{there is }\ell\geq j+m\text{ such that }X_{jt_{0}}F\cap X_{\ell t_{0}}F\neq\varnothing)$
@@
$\mathbb{P}_{\mu_{k}}\left(0\text{ is poorly }m_{1}\text{-stabilized}\right)\leq\mathbb{P}_{\mu_{k}}\left(0\text{ is poorly }m_{2}\text{-stabilized}\right).$
@@
This is a consequence of the fact that if $m_{1}\geq m_{2}$, then the event where 0 is poorly $m_{1}$-stabilized is contained in the event where 0 is poorly $m_{2}$-stabilized. ∎
@@
$\mathbb{P}_{\mu_{k}}\left(j\text{ is poorly }m\text{-stabilized}\right)$ $\leq\mathbb{P}_{\mu_{k}}\left(0\text{ is poorly }m\text{-stabilized}\right)$
@@
$\leq|F|\sum_{j=0}^{\left\lfloor n/t_{0}\right\rfloor-n_{0}}\mathbb{P}_{\mu_{k}}\left(j\text{ is poorly }n_{0}\text{-stabilized}\right)$
@@
Let $t_{0}\geq 1$ and consider any finite subset $F\subseteq B$ with $e_{B}\in F$ . Then for every $\varepsilon>0$, there exists $K,n_{0}\geq 1$ such that for all $k\geq K$ and $n>n_{0}t_{0}$ we have $H_{\mu_{k}}(\mathcal{U}_{n}(n_{0},t_{0},F)\mid\mathcal{P}_{n}^{t_{0}})<\varepsilon n.$
@@
[…]
@@
$\mathbb{E}_{\mu_{k}}(|\mathcal{V}_{n}(n_{0},t_{0},F)\backslash J_{n_{0}}|)\leq n_{0}|F|\mathbb{E}_{\mu_{k}}(|\mathcal{U}_{n}(n_{0},t_{0},F)|).$
@@
$H_{\mu_{k}}\Big(\mathcal{V}_{n}(n_{0},t_{0},F))\leq\left(\frac{n}{t_{0}}-n_{0}+1\right)\log\left(2\right)\leq\left(\frac{n}{T}\right)\log\left(2\right)=\varepsilon n,$